One ship is sailing south at a rate of 5 knots, and other is sailing east at a rate of 10 knots. At 2 pm, the second ship was at the place occupied by the first one hour before. At what time was the distance between the ships not changing?


1
Expert's answer
2021-01-14T17:53:38-0500

The distance between ships will be

"s(t)=\sqrt{(10t)^2+(5+5t)^2}"

here t=0 at 2pm.


"s'(t)=\frac{2*10t*10+2(5+5t)*5}{2\sqrt{(10t)^2+(5+5t)^2}}"


the distance was minimum when s'(t)=0, and in this moment

the distance between ships was not changing, so


"2*10t*10+2*(5+5t)*5=0"

"200t+50+50t=0"

"250t=-50"

"t=-\frac{1}{5}"


To find moment of time:

"2pm-\frac{1}{5}=1.48pm"


Answer: 1.48pm.